OpenStudy (anonymous):

Solve for x and y x^2-8i=(y-3)i+16

OpenStudy (anonymous):

x=+/-4 and y=-5?

OpenStudy (anonymous):

\[x^2 - 8i - yi + 3i = 16\] \[x^2 - i(5 + y) = 16 - 0i\] Compare for x and y...

OpenStudy (anonymous):

Yes you are right in solving this..

OpenStudy (anonymous):

but?